# PackKV: Reducing KV Cache Memory Footprint through LLM-Aware Lossy Compression

### v1

@@ I Introduction
Transformer-based large language models (LLMs) have revolutionized natural language processing, enabling breakthroughs in diverse tasks [brown2020language, taylor2022galactica]. The self-attention mechanism allows models to capture long-range dependencies and contextual information. However, these capabilities come at a significant computational and memory cost during inference with long input contexts, where the memory footprint of the key-value (KV) cache becomes a major bottleneck [kwon2023efficient, pope2023efficiently].
This growing footprint severely constrains the inference performance, limiting the achievable context length, reducing batch size, or impeding the deployment of LLMs on memory-constrained hardware [kwon2023efficient, liu2023scissorhands]. The KV cache stores intermediate key and value tensors for each token processed by the model and is reused during subsequent decoding steps to avoid redundant computation. As sequence length and batch size increase, the cache size grows linearly and can consume a substantial portion of GPU memory, sometimes exceeding the memory footprint of the model weights themselves [liu2023scissorhands]. For example, LLaMA2-30B inference with a context length of 32,000 and a batch size of 8 can produce over 100 GB of KV cache, surpassing the model size itself of 60 GB in float16.
To address these challenges, recent studies mainly leverage three approaches to reduce KV cache size: quantization, pruning, and GPU-CPU migration. Quantization-based methods [liu2024kivi, hooper2024kvquant], such as KIVI, aim to carefully design quantization strategies that minimize the impact on model accuracy. However, these techniques often achieve modest compression ratios unless combined with additional encoding, which introduces overhead and limits their applicability in latency-sensitive LLM inference. Pruning-based methods [wang2021spatten, zhang2024q], such as Q-Hitter, selectively discard KV pairs that are predicted to be unimportant for future decoding. While effective in some cases, these methods suffer from unpredictable attention callbacks, leading to either costly KV recomputation or substantial accuracy degradation. GPU-CPU migration is a traditional approach for handling memory overflows, offloading KV cache data to CPU memory [sheng2023flexgen]. Although this mitigates GPU memory pressure, it significantly degrades inference performance due to data transfer latency and complex scheduling overhead.
@@ II-A Large Language Model Inference
LLMs have become foundational in natural language processing, demonstrating remarkable capabilities in text generation [achiam2023gpt]. The inference process involves auto-regressive decoding, where the model generates tokens one by one, with each new token depending on previously generated ones and the initial prompt [radford2018improving, vaswani2017attention].
To enable efficient attention to preceding tokens without recomputing states at every generation step, Transformer models maintain a Key–Value (KV) cache [shazeer2019fast]. The core component enabling this process is the attention mechanism [vaswani2017attention]. During inference, Key (K) and Value (V) tensors are computed for each token within self-attention layers; these K and V tensors are then reused to process future tokens and collectively form the KV cache [shazeer2019fast]. The KV cache is organized as [context_len,head_num,head_dim][context\_len,head\_num,head\_dim], where context_lencontext\_len is the sequence length, head_numhead\_num is the number of attention heads, and head_dimhead\_dim is the head dimension.
@@
KV cache participates in two stages during inference [agrawal2023sarathi]. Prefill Stage: Initial prompt processing generates the initial KV cache (Figure 1(a)). Decode Stage: Auto-regressive token generation where new KV vectors are appended to the cache and the entire cache is used for attention computation via matrix-vector multiplications (Figure 1(b)).
By rapidly inflating GPU memory usage, the KV cache often becomes the bottleneck that constrains sequence length and batch size, degrading performance or even preventing inference on memory-constrained hardware. Although it enables efficient inference, the cache grows linearly with sequence length and batch size and can consume more GPU memory than model weights [kwon2023efficient, liu2023scissorhands]. Existing remedies include KV-cache quantization [liu2024kivi, hooper2024kvquant], which provides limited reduction at near-original accuracy, and GPU–CPU offloading [sheng2023flexgen], which lowers GPU memory needs at the cost of substantial performance overhead.
@@ II-B Data Reduction in LLMs
In large neural networks, particularly LLMs, lossy compression can be applied to large data structures like activations or caches to mitigate memory bottlenecks [han2016deep_compression, jacob2018quantization, banner2019post, frantar2022gptq, jin2019deepsz, jin2021comet]. For example, quantization is a form of lossy compression that reduces the precision of numerical data by mapping continuous or high-precision values (e.g., float32) to a smaller, discrete set of values such as integers. In fact, this inherently reduces the entropy of the data, making it smaller and often more amenable to lossless encoding. In image, video, and scientific lossy compression, entropy encoding and spatial encoding are commonly added to further reduce data size [wallace1992jpeg, taubman2002embedded, lindstrom2014fixed, ohm2012comparison, di2016fast, tao2017significantly, sz18]. However, the low throughput of these post-quantization encoders makes them impractical for real-world LLM inference, where latency is critical. For example, CacheGen [liu2024cachegen] integrates such encoders into the KV Cache quantization pipeline and achieves throughput below 1 GB/s, which is an improvement over network transmission, but insufficient compared to the GPU-CPU transmission and GPU memory bandwidth.
Another challenge in applying lossy compression is to manage the trade-off between the desired compression ratio and the potential impact on model accuracy. Applying quantization requires careful consideration of the data’s characteristics and the specific downstream task. Different quantization strategies, such as varying granularity (e.g., token/channel/block-wise) or using adaptive error bounds, can be employed to control the information loss and minimize accuracy degradation [liu2024kivi].
@@ II-C GPU architecture
The GPU memory hierarchy plays a crucial role in performance optimization [mei2016dissecting]. Global memory provides the largest capacity with high latency, making efficient access patterns essential. For example, coalesced memory access can significantly improve bandwidth utilization. While shared memory in each SM is a programmer-managed cache shared among threads within a block.
@@ III-B2 Quantization
Several quantization granularities can be applied to the KV cache tensor (Figure 5). KIVI [liu2024kivi] uses fine-grained channel-wise quantization for K cache and token-wise for V cache, exploiting channel correlations (i.e., column-wise in Figure 4). However, quantization-only methods have limited flexibility as bit-widths must be integers (2, 3, or 4 bits). Higher compression ratios require reducing bit-width, but this demands smaller quantization units to maintain model accuracy, increasing metadata overhead. For KIVI’s 2-bit K quantization with 64-element blocks which is recommended by KIVI (even 128 is not recommended), the actual compression ratio is only 6.4 (not 8) due to FP16 metadata. If bit-width falls back to 3 bits, compression drops to 4.57 which is close to that of coarse-grained 4-bit quantization. This reveals quantization’s dilemma: Using finer quantization granularity to reach lower bit widths inflates metadata, capping overall compression.
@@ III-B3 Encode-aware Repacking and Bit-packing Encoding
Problem Analysis: The formalized problem reduces to an optimization version of the set partition problem, which is NP-hard [rasmussen2011optimisation]. Given this complexity, we propose a greedy algorithm 1 that incrementally constructs packs by selecting vectors that locally minimize incremental cost:
@@ IV-A Experiment Setup
Our evaluation metrics are Model Accuracy, Compression Ratio, and Decompression Throughput (including computation). We selected six models (i.e., Llama2-7B/13B, Llama3.1-8B, DeepSeek-R1-Llama-8B, Ministral-8B-2410, Phi-4) and six benchmarks (i.e., CoQA, GSM8K, MMLU, Winogrande, GPQA (Diamand[zeroshot, n_shot] as GPQA_D), SQuAD Completion as SQuAD_C) for accuracy evaluation. We use KIVI quantization [liu2024kivi] as the baseline for accuracy and compression ratio. For decompression throughput, we compare our computation-aware decompression+matrix_vector_multiplication kernel with original cuBLAS matrix vector multiplication.

### v2

@@ I Introduction
Transformer-based large language models (LLMs) have revolutionized natural language processing, enabling breakthroughs in diverse tasks [4, 27]. The self-attention mechanism allows models to capture long-range dependencies and contextual information. However, these capabilities come at a significant computational and memory cost during inference with long input contexts, where the memory footprint of the key-value (KV) cache becomes a major bottleneck [12, 20].
This growing footprint severely constrains the inference performance, limiting the achievable context length, reducing batch size, or impeding the deployment of LLMs on memory-constrained hardware [12, 16]. The KV cache stores intermediate key and value tensors for each token processed by the model and is reused during subsequent decoding steps to avoid redundant computation. As sequence length and batch size increase, the cache size grows linearly and can consume a substantial portion of GPU memory, sometimes exceeding the memory footprint of the model weights themselves [16]. For example, LLaMA2-30B inference with a context length of 32,000 and a batch size of 8 can produce over 100 GB of KV cache, surpassing the model size itself of 60 GB in float16.
To address these challenges, recent studies mainly leverage three approaches to reduce KV cache size: quantization, pruning, and GPU-CPU migration. Quantization-based methods [17, 8], such as KIVI, aim to carefully design quantization strategies that minimize the impact on model accuracy. However, these techniques often achieve modest compression ratios unless combined with additional encoding, which introduces overhead and limits their applicability in latency-sensitive LLM inference. Pruning-based methods [30, 31], such as Q-Hitter, selectively discard KV pairs that are predicted to be unimportant for future decoding. While effective in some cases, these methods suffer from unpredictable attention callbacks, leading to either costly KV recomputation or substantial accuracy degradation. GPU-CPU migration is a traditional approach for handling memory overflows, offloading KV cache data to CPU memory [24]. Although this mitigates GPU memory pressure, it significantly degrades inference performance due to data transfer latency and complex scheduling overhead.
@@ II-A Large Language Model Inference
LLMs have become foundational in natural language processing, demonstrating remarkable capabilities in text generation [1]. The inference process involves auto-regressive decoding, where the model generates tokens one by one, with each new token depending on previously generated ones and the initial prompt [21, 28].
To enable efficient attention to preceding tokens without recomputing states at every generation step, Transformer models maintain a Key–Value (KV) cache [23]. The core component enabling this process is the attention mechanism [28]. During inference, Key (K) and Value (V) tensors are computed for each token within self-attention layers; these K and V tensors are then reused to process future tokens and collectively form the KV cache [23]. The KV cache is organized as [context_len,head_num,head_dim][context\_len,head\_num,head\_dim], where context_lencontext\_len is the sequence length, head_numhead\_num is the number of attention heads, and head_dimhead\_dim is the head dimension.
@@
KV cache participates in two stages during inference [2]. Prefill Stage: Initial prompt processing generates the initial KV cache (Figure 1(a)). Decode Stage: Auto-regressive token generation where new KV vectors are appended to the cache and the entire cache is used for attention computation via matrix-vector multiplications (Figure 1(b)).
By rapidly inflating GPU memory usage, the KV cache often becomes the bottleneck that constrains sequence length and batch size, degrading performance or even preventing inference on memory-constrained hardware. Although it enables efficient inference, the cache grows linearly with sequence length and batch size and can consume more GPU memory than model weights [12, 16]. Existing remedies include KV-cache quantization [17, 8], which provides limited reduction at near-original accuracy, and GPU–CPU offloading [24], which lowers GPU memory needs at the cost of substantial performance overhead.
@@ II-B Data Reduction in LLMs
In large neural networks, particularly LLMs, lossy compression can be applied to large data structures like activations or caches to mitigate memory bottlenecks [7, 9, 3, 6, 10, 11]. For example, quantization is a form of lossy compression that reduces the precision of numerical data by mapping continuous or high-precision values (e.g., float32) to a smaller, discrete set of values such as integers. In fact, this inherently reduces the entropy of the data, making it smaller and often more amenable to lossless encoding. In image, video, and scientific lossy compression, entropy encoding and spatial encoding are commonly added to further reduce data size [29, 26, 14, 19, 5, 25, 13]. However, the low throughput of these post-quantization encoders makes them impractical for real-world LLM inference, where latency is critical. For example, CacheGen [15] integrates such encoders into the KV Cache quantization pipeline and achieves throughput below 1 GB/s, which is an improvement over network transmission, but insufficient compared to the GPU-CPU transmission and GPU memory bandwidth.
Another challenge in applying lossy compression is to manage the trade-off between the desired compression ratio and the potential impact on model accuracy. Applying quantization requires careful consideration of the data’s characteristics and the specific downstream task. Different quantization strategies, such as varying granularity (e.g., token/channel/block-wise) or using adaptive error bounds, can be employed to control the information loss and minimize accuracy degradation [17].
@@ II-C GPU architecture
The GPU memory hierarchy plays a crucial role in performance optimization [18]. Global memory provides the largest capacity with high latency, making efficient access patterns essential. For example, coalesced memory access can significantly improve bandwidth utilization. While shared memory in each SM is a programmer-managed cache shared among threads within a block.
@@ III-B2 Quantization
Several quantization granularities can be applied to the KV cache tensor (Figure 5). KIVI [17] uses fine-grained channel-wise quantization for K cache and token-wise for V cache, exploiting channel correlations (i.e., column-wise in Figure 4). However, quantization-only methods have limited flexibility as bit-widths must be integers (2, 3, or 4 bits). Higher compression ratios require reducing bit-width, but this demands smaller quantization units to maintain model accuracy, increasing metadata overhead. For KIVI’s 2-bit K quantization with 64-element blocks which is recommended by KIVI (even 128 is not recommended), the actual compression ratio is only 6.4 (not 8) due to FP16 metadata. If bit-width falls back to 3 bits, compression drops to 4.57 which is close to that of coarse-grained 4-bit quantization. This reveals quantization’s dilemma: Using finer quantization granularity to reach lower bit widths inflates metadata, capping overall compression.
@@ III-B3 Encode-aware Repacking and Bit-packing Encoding
Problem Analysis: The formalized problem reduces to an optimization version of the set partition problem, which is NP-hard [22]. Given this complexity, we propose a greedy algorithm 1 that incrementally constructs packs by selecting vectors that locally minimize incremental cost:
@@ IV-A Experiment Setup
Our evaluation metrics are Model Accuracy, Compression Ratio, and Decompression Throughput (including computation). We selected six models (i.e., Llama2-7B/13B, Llama3.1-8B, DeepSeek-R1-Llama-8B, Ministral-8B-2410, Phi-4) and six benchmarks (i.e., CoQA, GSM8K, MMLU, Winogrande, GPQA (Diamand[zeroshot+n_shot] as GPQA_D), SQuAD Completion as SQuAD_C) for accuracy evaluation. We use KIVI quantization [17] as the baseline for accuracy and compression ratio. For decompression throughput, we compare our computation-aware decompression+matrix_vector_multiplication kernel with original cuBLAS matrix vector multiplication.